OpenStudy (anonymous):

Jacob invested $300 in a savings account and earned $81 in interest at the end of 9 years. What was the interest rate

OpenStudy (anonymous):

1529

OpenStudy (hanner_b_nanner):

$81/9= $9 per year. Make a proportion... 9/300=x/100 9x100=900/300=3 3% interest rate! :))